# DiffMesh: A Motion-aware Diffusion Framework for Human Mesh Recovery from Videos

### v6

@@ Abstract
Figure 1: Different approaches of applying diffusion model for video-based HMR, the input frame $f$ is 3 for simplicity and the number of steps is $N$. Here $x_{i}$ and $c_{i}$ denote the mesh and conditional features of $i_{th}$ frame, respectively. (a) Baseline 1: The diffusion model is applied for each frame individually. The total steps are $f\times N$, which is computationally expensive. Moreover, motion is ignored during the denoising process, leading to non-smooth motion predictions. (b) Baseline 2: The features from ground truth mesh feature of each frame $x_{i}$ are concatenated to unified features $x_{unified}$ during the forward process. To obtain the mesh of each frame, features are split after the denoising process. Although this strategy reduces the total steps to $N$, it doesn’t effectively model human motion during denoising. (c) Our proposed DiffMesh: We consider the inherent motion patterns within the forward process and the reverse process when utilizing the diffusion model, leading to smooth and accurate motion predictions. The total number of steps remains as $N$. For more detailed information, please refer to Sec 3.3.
@@ 3.1 Preliminary
$\small p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\hat{\beta_{t}}}\cdot x_{0},(1-\hat{\beta_{t}})\cdot\textbf{I}),\vspace{-10pt}$ (2)
@@
$\small x_{t}=\sqrt{\hat{\beta_{t}}}\cdot x_{0}+\sqrt{(1-\hat{\beta_{t}})}\cdot z_{t},\vspace{-10pt}$ (3)
@@
$\small x_{t-1}=\frac{1}{\sqrt{\beta_{t}}}(x_{t}-\frac{1-\beta_{t}}{\sqrt{1-\hat{\beta_{t}}}}\epsilon_{\theta}(x_{t},t))+\sigma_{t}z,\vspace{-5pt}$ (5)
@@ 3.2 Diffusion-based Baselines
Figure 5: The architecture of DiffMesh: Our framework takes input sequence with $f$ frames, with the objective of outputting a human mesh sequence consisting of $f$ frames. We model the forward human motion across frames similar to the mechanism of introducing noise in the forward process. We assume that the human motion sequence will eventually reach the initial distribution within total $N$ steps. (additional $N-f+1$ steps are necessary from $x_{f}$ to reach the initial distribution $x_{end}$). Consequently, we utilize a transformer-based diffusion model to sequentially produce the decoded features during the reverse process. The final human mesh sequence is returned by a mesh head using SMPL [34] human body model. The structure of DN (diffusion network) is illustrated in Fig. 6.
@@ 3.3 DiffMesh
$\small x_{i+1}=\sqrt{\beta_{i}}\cdot x_{i}+\sqrt{(1-\beta_{i})}\cdot{\color[rgb]{1,.5,0}m_{i}},\vspace{-5pt}$ (6)
where $m_{i}$ represents the motion pattern from $i_{th}$ frame to $(i+1)_{th}$ frame. Consequently, the mesh feature of the first frame $x_{1}$ can be gradually diffused into the mesh feature of the last frame $x_{f}$ after $(f-1)$ forward diffusion steps. Importantly, achieving the last frame $x_{f}$ does not denote the conclusion of the forward diffusion process. We assume that mesh feature with adding human motion will eventually reach the initial distribution $x_{end}$ through total $N$ diffusion steps (similar to the total $N$ steps to get the Gaussian distribution in the conventional diffusion model). More details such as mathematical proof and ablation study are provided in the supplementary Sec. 3. To arrive at this initial distribution $x_{end}$, DiffMesh requires $N-f+1$ additional forward diffusion steps from $x_{f}$. Therefore, the total number of steps remains $N$.
Reverse Process: Our objective during the reverse process is to obtain output features $x^{\prime}_{1}\cdots x^{\prime}_{f}$. Initiating from the initial distribution $x_{end}$, we design a transformer-based diffusion model, as depicted in Fig. 5, to sequentially produce decoded features. After $N-f+1$ steps, our diffusion model yields $x^{\prime}_{f}$, which serves as the target feature for the final frame $f$. Iteratively, we generate the output features $x^{\prime}_{i}$ where $i\in\left\{f-1,\dots 1\right\}$ over $f-1$ steps. The total number of steps remains $N$. In contrast to Baseline 2, DiffMesh demonstrates the capability to decode specific motion patterns within the same number of steps, while generating more precise and smooth output (verified in Sec. 4.4).
@@ B.4 More Details about the Architecture
Conditional features generation block: Our chosen backbone to extract features for both our proposed method and the baselines is ResNet-50 [14] or DSTformer [72]. After extracting features from each frame $b_{i}$, where $i\in\left\{1,\dots f\right\}$, using the backbone, our goal is to generate $N$ conditional features to be utilized during the reverse process. To achieve this, we pad additional $N-f$ zero features, $b_{f+1},\cdots,b_{N}$. Then, we combine them with the existing features, creating new features $b\in\mathbb{R}^{N\times D}$, where $D$ represents the embedded dimension. Subsequently, we apply a transformer block [9] to model these features and return the required conditional features denoted as $c\in\mathbb{R}^{N\times D}$.
@@ Appendix D Human Mesh Visualization
In our experimental setup, we utilize 16 input frames, and the total number of steps is set to 30. In the reverse motion process, DiffMesh outputs [$y_{1},y_{2}\cdots,y_{30}$] over 30 steps. For the output mesh sequence of 16 frames, we use [$y_{1},y_{2}\cdots,y_{16}$]. Additionally, we generate the mesh from [$y_{16},y_{17}\cdots,y_{30}$], as visually depicted in Fig 12. This visualization illustrates the trend of the generated human mesh gradually decoding toward the desired human mesh of each input frame.

### v7

@@ Abstract
Figure 1: Different approaches of applying diffusion model for video-based HMR, the input frame $f$ is 3 for simplicity and the number of steps is NN. Here $x_{i}$ and $c_{i}$ denote the mesh and conditional features of $i_{th}$ frame, respectively. (a) Baseline 1: The diffusion model is applied for each frame individually. The total steps are $f\times N$, which is computationally expensive. Moreover, motion is ignored during the denoising process, leading to non-smooth motion predictions. (b) Baseline 2: The features from ground truth mesh feature of each frame $x_{i}$ are concatenated to unified features $x_{unified}$ during the forward process. To obtain the mesh of each frame, features are split after the denoising process. Although this strategy reduces the total steps to $N$, it doesn’t effectively model human motion during denoising. (c) Our proposed DiffMesh: We consider the inherent motion patterns within the forward process and the reverse process when utilizing the diffusion model, leading to smooth and accurate motion predictions. The total number of steps remains as $N$. For more detailed information, please refer to Sec $3.$
@@ 3.1 Preliminary
$\small p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\hat{\beta_{t}}}\cdot x_{0},(1-\hat{\beta_{t}})\cdot\textbf{I}),\vskip-10.0pt$ (2)
@@
$\small x_{t}=\sqrt{\hat{\beta_{t}}}\cdot x_{0}+\sqrt{(1-\hat{\beta_{t}})}\cdot z_{t},\vskip-10.0pt$ (3)
@@
$\small x_{t-1}=\frac{1}{\sqrt{\beta_{t}}}(x_{t}-\frac{1-\beta_{t}}{\sqrt{1-\hat{\beta_{t}}}}\epsilon_{\theta}(x_{t},t))+\sigma_{t}z,\vskip-5.0pt$ (5)
@@ 3.2 Diffusion-based Baselines
Figure 5: The architecture of DiffMesh: Our framework takes input sequence with $f$ frames, with the objective of outputting a human mesh sequence consisting of $f$ frames. We model the forward human motion across frames similar to the mechanism of introducing noise in the forward process. We assume that the human motion sequence will eventually reach the initial distribution within total NN steps. (additional Nf+1N-f+1 steps are necessary from $x_{f}$ to reach the initial distribution $x_{end}$). Consequently, we utilize a transformer-based diffusion model to sequentially produce the decoded features during the reverse process. The final human mesh sequence is returned by a mesh head using SMPL [34] human body model. The structure of DN (diffusion network) is illustrated in Fig. 6.
@@ 3.3 DiffMesh
$\small x_{i+1}=\sqrt{\beta_{i}}\cdot x_{i}+\sqrt{(1-\beta_{i})}\cdot{\color[rgb]{1,.5,0}m_{i}},\vskip-5.0pt$ (6)
where $m_{i}$ represents the motion pattern from $i_{th}$ frame to $(i+1)_{th}$ frame. Consequently, the mesh feature of the first frame $x_{1}$ can be gradually diffused into the mesh feature of the last frame $x_{f}$ after (f1)(f-1) forward diffusion steps. Importantly, achieving the last frame $x_{f}$ does not denote the conclusion of the forward diffusion process. We assume that mesh feature with adding human motion will eventually reach the initial distribution $x_{end}$ through total $N$ diffusion steps (similar to the total $N$ steps to get the Gaussian distribution in the conventional diffusion model). More details such as mathematical proof and ablation study are provided in the supplementary Sec. 3. To arrive at this initial distribution $x_{end}$, DiffMesh requires Nf+1N-f+1 additional forward diffusion steps from $x_{f}$. Therefore, the total number of steps remains $N$.
Reverse Process: Our objective during the reverse process is to obtain output features $x^{\prime}_{1}\cdots x^{\prime}_{f}$. Initiating from the initial distribution $x_{end}$, we design a transformer-based diffusion model, as depicted in Fig. 5, to sequentially produce decoded features. After Nf+1N-f+1 steps, our diffusion model yields $x^{\prime}_{f}$, which serves as the target feature for the final frame $f$. Iteratively, we generate the output features $x^{\prime}_{i}$ where $i\in\left\{f-1,\dots 1\right\}$ over f1f-1 steps. The total number of steps remains $N$. In contrast to Baseline 2, DiffMesh demonstrates the capability to decode specific motion patterns within the same number of steps, while generating more precise and smooth output (verified in Sec. 4.4).
@@ B.4 More Details about the Architecture
Conditional features generation block: Our chosen backbone to extract features for both our proposed method and the baselines is ResNet-50 [14] or DSTformer [72]. After extracting features from each frame $b_{i}$, where $i\in\left\{1,\dots f\right\}$, using the backbone, our goal is to generate $N$ conditional features to be utilized during the reverse process. To achieve this, we pad additional NfN-f zero features, $b_{f+1},\cdots,b_{N}$. Then, we combine them with the existing features, creating new features $b\in\mathbb{R}^{N\times D}$, where $D$ represents the embedded dimension. Subsequently, we apply a transformer block [9] to model these features and return the required conditional features denoted as $c\in\mathbb{R}^{N\times D}$.
@@ Appendix D Human Mesh Visualization
In our experimental setup, we utilize 16 input frames, and the total number of steps is set to 30. In the reverse motion process, DiffMesh outputs $y_{1},y_{2}\cdots,y_{30}$] over 30 steps. For the output mesh sequence of 16 frames, we use $y_{1},y_{2}\cdots,y_{16}$]. Additionally, we generate the mesh from $y_{16},y_{17}\cdots,y_{30}$], as visually depicted in Fig 12. This visualization illustrates the trend of the generated human mesh gradually decoding toward the desired human mesh of each input frame.